A net force of 200 N is applied to a large object, causing its velocity to change from 30km/h to 20 km/h in 2.3 seconds. What is the objects acceleration? What is its mass?

1
Expert's answer
2022-04-12T16:13:02-0400

The object acceleration

"a=\frac{\Delta v}{\Delta t}=\frac{(30-20)\div 3.6}{2.3}=1.2\;\rm m/s^2"

The mass of the object

"m=\frac{F}{a}=\frac{200}{1.2}=166\:\rm kg"